The market demand and supply functions for an electronic typewriter are

QD = 20,000 – 10P

QS = 10,000 + 20P

               Western Electronics Inc. is one of many firms in this perfectly competitive industry. Its marginal and average cost functions are

                                               MC = 10 + 2Q

                                               AC = (1,000/Q) + 10 + Q

               a/ What is the profit-maximizing output rate for Western?

               b/ How much economic profit will be earned at that rate of output?


1
Expert's answer
2022-03-21T12:49:29-0400

A) Given that the demand function is

"Q_D=20000-10P"

And "MC=10+2Q"

"TR=P\u00d7Q"

"P=2000-0.1Q"

"TR=(2000-0.1Q)Q"

"TR=2000Q-0.1Q^2"

The profit maximizing output is given at the equilibrium, "MR=MC"

Where MR is:

"MR=TR'=2000-0.2Q"

"2000-0.2Q=10+2Q"

"Q^*=905"


B) "\pi=TR-TC" at Q=905

Where TC is

"TC=AC\u00d7Q"

"TC=(\frac{1000}{Q}+10+Q)Q"

"TC=1000+10Q+Q^2"

"\therefore \pi=2000Q-0.1Q^2-(1000+10Q+Q^2)"

"2000(905)-0.1(905)^2-(1000+10(905)+905^2"

"\pi=899022.5"